OpenStudy (anonymous):

Find the 25th term of the following sequence: 20, 18, 16... Show all work a^n = a^1 + (n - 1)d

OpenStudy (anonymous):

@kirbykirby @vlltimelow @amistre64 @abb0t @quickstudent @ganeshie8 Please help me out with this. I also have a couple more questions.

OpenStudy (amistre64):

is that a subn, or a to the nth power?

OpenStudy (anonymous):

subn, sorry I didn't clarify that.

OpenStudy (amistre64):

since a1 is given, and n is given, all thats required is to find d

OpenStudy (amistre64):

20+d = 18 18+d = 16 etc ...

OpenStudy (kirbykirby):

\(a_{25}=a_1+(25-1)d\) Here, \(a_1\) is the first term, 20 Now, \(d\) is the common difference between consecutive terms... i.e. 18 - 20 = -2 = d 16 - 18 = -2 = d

OpenStudy (anonymous):

That's what the formula looks like by the way

OpenStudy (anonymous):

So what would I do now?

OpenStudy (amistre64):

determine the value of the 25th term of course

OpenStudy (kirbykirby):

replace \(a_1\) = 20, d = -2 into the formula and you get \(a_{25}\)

OpenStudy (anonymous):

So a25 = 20 + (25 - 1)-2

OpenStudy (kirbykirby):

um "yes" although the -2 should be written in parentheses... (25 - 1)(-2) since you are multiplying it